# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ 3      $A\xleftarrow{}AddtoArchive(samples,MS,FDR);$
Algorithm 2 describes the third step of TEASMA. As illustrated in Figure 2, this step is performed on a set of unlabeled test inputs before deployment. Given a test set $T$ for the DNN $N$ to assess, TEASMA starts by loading the same set of mutants $M$ generated in the first step (line 1) and the prediction model $PM$ built in the second step (line 2). To predict the FDR of a test set $T$, TEASMA first calculates MS (line 3) using the same MS calculation approach $MSFormula$ that has been used in the first step. Subsequently, the MS of the test set $T$ is used to predict its FDR based on the loaded prediction model $PM$ (line 4). TEASMA provides engineers not only with the predicted FDR of $T$ but also with its prediction interval $PI$ to enable them to decide on the adequacy of a test set considering the lower or higher bounds of the PI (Section III-E).
We argue that our proposed solution—assuming an accurate FDR prediction model can be built specifically for each DNN using its training set—is practical. Indeed, TEASMA relies on post-training MOs which are directly applied to the DNN model and do not entail any re-training, thus avoiding the main challenge with pre-training MOs. Algorithm 1 is executed once for a specific DNN to generate mutants and create the prediction model. Subsequently, Algorithm 2 can be employed multiple times to assess many test sets without incurring additional costs. In practice, such assessment helps guide the selection of test inputs and significantly reduces test costs by only labeling a test set once it is considered adequate. However, since we build the FDR prediction model using the training set and use it to determine the adequacy of a test set, we implicitly assume that the two sets have similar distributions. This is, however, a common assumption for the training of DNNs to be considered adequate.
@@
III-A Calculating Mutation Scores
To calculate MS precisely, in addition to configuring MOs correctly, we need to complement the mutant generation process by identifying and excluding mutants that are trivial or equivalent. TEASMA uses DeepMutation++ [12], the most recent SoTA mutation testing tool for DNNs that implements eight post-training MOs suitable for FNNs. Some of the implemented MOs need to be configured by setting values for their parameters. For example, the Neuron Effect Blocking (NEB) operator is configured by determining the number of neurons it affects. By default, the DeepMutation++ tool configures this operator to sample a predefined ratio of 1% of the total neurons in a DNN model and limits the total number of generated mutants to 50. TEASMA uses the default configuration for generating mutants. Furthermore, it performs the same procedure proposed by the DeepMutation++ tool and uses its default thresholds to filter out some of the generated mutants. Specifically, it excludes mutants when they yield less than 90% of the original model’s accuracy or mispredict more than 20% of the inputs correctly predicted by the original model, as reported in the original papers [11, 12]. Additionally, TEASMA performs an extra step to filter out equivalent mutants. The identification of equivalent mutants is achieved through a procedure similar to that of DeepCrime [13], involving the execution of the entire training set against the mutants. Mutants that cannot be killed by any training input, are considered equivalent and hence excluded. TEASMA can employ various approaches for MS calculation as explained in Section II-B. Consequently, in our experiment, we examine the correlation between FDR and MS calculated using each approach and empirically determine which approach results in the most accurate prediction model.
@@ III-B Estimating Faults
TEASMA estimates faults in the DNN model using the approach described in Section II-E based on the mispredicted inputs of the training set. As we mentioned earlier, we implicitly assume that both the training set and the test set have a similar distribution. Except for transfer learning [28] where there is an expected distribution shift between the training set and the test set, this default assumption is correct and necessary for the training process to be considered adequate. Therefore, we can use the training set of the DNN model to identify faults of the DNN model and create its specific FDR prediction model. Note that, unlike test inputs, training inputs are labeled, and mispredicted inputs can be identified. To accomplish this, the entire training set is executed against the original DNN model, the mispredicted inputs are identified, and subsequently, the fault identification approach [6] is applied to these mispredicted inputs. Note that TEASMA builds a specific fault prediction model for each individual DNN based on its training set. Consequently, the performance of TEASMA can potentially be enhanced by integrating new fault identification approaches and creating more precise prediction models.
@@ III-C Sampling Input Subsets
TEASMA, selects input subsets from the training set, calculates the MS and FDR of each subset, and builds a regression model. For this purpose, it requires a large number of input subsets that possess diverse MS and FDR values. Therefore, starting with sampling input subsets with a small size, TEASMA updates the sampling size and continues with larger or smaller sizes where needed, until the entire spectrum of FDR with minimum and maximum values ($0\leq FDR\leq 1$) is covered. To build an accurate prediction model, TEASMA requires the sampling of multiple subsets along the size range. The number of subsets sampled for each size is determined through an input parameter ($SN$) provided to TEASMA, which we set to 300 in our experiments. Sampling is performed with replacement, allowing for the potential inclusion of the same test input in multiple subset samples. This approach serves the purpose of including diverse testing inputs and variations of MS and FDR values across various samples. Since one of the MS formulas we use (Equation (2)) considers the number of images from each class $C_{i}$ in an input subset, TEASMA complements the conventional approach of creating randomly sampled subsets by creating a set of uniformly sampled subsets. Random subsets may contain any number of input images from each class $C_{i}$ but uniform subsets contain an equal number for each class $C_{i}$. Uniform subsets may achieve higher MS values than random subsets based on Equation (2). Therefore, we perform our experiments with both types of input subsets and empirically determine which one serves our purpose better.
@@ III-D Regression Analysis
In this step, TEASMA uses all the sampled input subsets from the training set to perform regression analysis between MS and FDR. Given that post-training MOs may not be representative of real faults, we can expect non-linear relationships. In fact, as our subsequent results demonstrate, the shape of the relationship between MS and FDR is often highly non-linear and significantly varies across DNNs. Therefore, TEASMA considers quadratic and exponential regression models, as well as regression trees when no regression function fits well. Regression trees enable the optimal partition of the MS range into optimal sub-ranges with similar FDR values while avoiding overfitting. Note that the best model shapes may differ across DNNs because of differences in their architecture and training sets. Consequently, for each DNN model, the best shape for the regression model is determined empirically, prioritizing simpler linear models when they fare well. Regardless of its shape, the regression model must be sufficiently accurate to enable practically useful FDR predictions. To measure the goodness of fit for each regression model, TEASMA calculates the resulting coefficient of determination ($R^{2}$). TEASMA also computes two common measures for evaluating the accuracy of predictions, namely the Mean Magnitude of Relative Error (MMRE) and the Root Mean Square Error (RMSE) to assess how accurately FDR can be predicted. To obtain more realistic results, TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average $R^{2}$, MMRE, and RMSE across all folds. Based on these metrics, engineers can determine for each DNN, the best regression model for predicting FDR based on MS.
@@ III-E Assessing Test Sets
In the last step, when provided with a test set, TEASMA computes its MS using the original DNN and the previously generated mutants. Subsequently, it predicts the test set’s FDR ($\widehat{FDR}$), including PIs based on the prediction model obtained from the previous steps. Given a certain confidence level (e.g., 95% in our experiments), such intervals specify the range of potential actual FDR values for test sets achieving a specific MS value and thus provide engineers with valuable insights into the uncertainty of FDR predictions. Depending on the context, one can then decide how conservative should decision-making be by considering the lower or higher bounds of the PI. The $\widehat{FDR}$ of a test set, if deemed sufficiently accurate based on its PI, can then be used to assess its capability to detect faults and therefore whether, in a given context, it is adequate to validate the DNN.
@@ IV Experimental Procedure
In this section, we describe the empirical evaluation of TEASMA, including the research questions we address, the subjects on which we conducted our evaluation and a comprehensive description of our experiments.
@@ IV-A Research Questions
Our experiments are designed to answer the following research questions:
RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
In this question, we investigate the output of the second step of TEASMA in Section III. We select a large number of input subsets from the training set, calculate their MS and FDR using the faults estimated in the original model as described in Section III-B, and investigate, for mutants derived from post-training MOs, if the MS of an input subset is significantly correlated with its FDR. In that case, we resort to regression analysis to model the relationship between MS and FDR and investigate if an accurate regression model can be built to predict a subset’s FDR from its MS. If that is the case, the predicted FDR can serve as a good basis on which to decide whether a test set provides sufficient confidence about the validation of the DNN before its deployment. Though such questions have been investigated for traditional software [15, 8], that has not been the case for DNNs.
RQ2: Can a test set’s FDR be accurately predicted from its MS using the regression model built on the training set?
In this question, we evaluate the last step of TEASMA in Section III and investigate if we can accurately predict a test set’s FDR using the regression model built in the previous step on the training set. For this purpose, we select a large number of input subsets from the test set and calculate their MS and FDR using the faults identified in the original model. Then, we use the regression model to predict FDR for each subset and compare it with the actual one, to assess whether, as expected, they have a linear relationship with a high correlation and a slope close to 1.
@@ IV-B Subjects
TABLE II: Information about experimental subjects
ID Input Training Test Model Epochs Train Test Number of mispredictions Number of faults
set set set accuracy accuracy (training) (training)
S1 MNIST 60,000 10,000 LeNet5 12 96% 96% 2218 307
S2 Cifar-10 50,000 10,000 Conv-8 20 94% 85% 2507 354
S3 ResNet20 100 97% 89% 1437 191
S4 Fashion-MNIST 60,000 10,000 LeNet4 20 94% 90% 3579 512
S5 SVHN 73257 26032 LeNet5 20 94% 87% 4257 607
S6 VGG16 10 97% 93% 5469 590
@@
We performed our study on different combinations of input sets and models. Since we rely on a SoTA fault estimation approach for DNNs that is tailored for image inputs, we selected a set of widely-used publicly available image input sets including MNIST [29], Fashion-MNIST [29], Cifar-10 [30], and SVHN [31]. Some of these input sets have also been used for evaluating SoTA mutation testing tools and operators [11, 10, 12, 13]. MNIST is a set of images of handwritten digits and Fashion-MNIST contains images of clothes that are associated with fashion and clothing items. Each one of MNIST and Fashion-MNIST represents a 10-class classification problem and consists of 70,000 grayscale images of a standardized 28x28 size. Cifar-10 is another widely-used input set that includes a collection of images from 10 different classes (e.g., cats, dogs, airplanes, cars). SVHN [31], a set of real-world house numbers collected by Google Street View, can be considered similar to MNIST since it includes images representing a sequence of digits from 10 classes of digits. The Cifar-10 and SVHN input sets contain 32x32 cropped colored images.
@@
In our experiments, we trained five SoTA DNN models using the above input sets: LeNet5, LeNet4, ResNet20, VGG16, and an 8-layer Convolutional Neural Network (Conv-8). Since not all of the post-training MOs are applicable to all DNN models, as described in Section II-A, we selected DNN models with different internal architectures. All the feasible combinations of DNN models and input sets we experiment with are referred to as subjects. Overall, we perform our experiment on six subjects including a wide range of diverse image inputs and DNN architectures. Table II lists the details of each subject including the size of test and training datasets, the number of epochs used to train the model, the accuracy of the DNN model on the training and test input sets, and the number of mispredicted inputs and distinct faults identified based on the training set. Note that we reused the training and test sets as defined in the original sources [29, 31, 30].
IV-C Experiments

### v3

@@ 5 $PM\xleftarrow{}LoadPredictionModel(N)$
Algorithm 2 describes the third step of TEASMA. As illustrated in Figure 2, this step is performed on a set of unlabeled test inputs. Given a test set $T$ for the DNN $N$ to assess, TEASMA starts by selecting the same configuration (e.g., generated mutants, MS calculation approach, or SC configurations) as in the first step (line 1) and the prediction model $PM$ built in the second step (line 2). To predict the FDR of a test set $T$, TEASMA first calculates AS (line 3), which is used in input of the prediction model $PM$ (line 4). TEASMA provides engineers not only with the predicted FDR of $T$ but also with its prediction interval $PI$ to enable them to decide on the adequacy of a test set considering the lower or higher confidence bounds of the prediction (Section III-E).
We argue that our proposed methodology—assuming an accurate FDR prediction model can be built specifically for each DNN using its training set—is practical. Indeed, even for the expensive MS calculation, TEASMA relies on post-training MOs, which are directly applied to the DNN model and do not entail any re-training, thus avoiding the main challenge with pre-training MOs. Algorithm 1 is executed once for a specific DNN to generate mutants and create the prediction model. Subsequently, Algorithm 2 can be employed multiple times to assess many test sets without incurring additional costs. In practice, such assessment helps guide the selection of test inputs and significantly reduces test costs by only labeling a test set once it is considered adequate. However, since we build the FDR prediction model using the training set and use it to determine the adequacy of a test set, we implicitly assume that the two sets have similar distributions. This is, however, a common assumption for the training of DNNs to be considered adequate. In cases where this assumption would not hold such that test results would show poor model accuracy, accurate test adequacy assessment is not necessary or useful, as the DNN obviously needs to be retrained to achieve any confidence.
@@
III-A Calculating AS
It is essential to note that TEASMA can employ any test set adequacy metric. However, the same metric must be used when calculating AS for both labeled training subsets (Step 1) and unlabeled test subsets (Step 3). In our study, the AS is calculated using four specific metrics: MS, DSC, LSC, and IDC. The MS calculation involves generating mutants and filtering out the trivial or equivalent ones. For SC calculation, parameters such as the lower and upper SA values are configured. To compute IDC, a VAE is trained using the training set. In Section IV-C, we describe the details of alternative AS calculations for each adequacy metric
@@ III-B Estimating Faults
It is essential to note that TEASMA can utilize any adequate fault identification approach. However, for the reasons outlined in Section II-D, this study utilizes the SOTA clustering approach introduced by Aghababaeyan et al. [18]. TEASMA estimates faults in the DNN model based on the mispredicted inputs of the training set. As we mentioned earlier, we implicitly assume that both the training set and the test set have a similar distribution. Except for transfer learning [35] where there is an expected distribution shift between the training set and the test set, this default assumption is necessary for the training process to be considered adequate. Therefore, we can use the training set of the DNN model to identify faults in the DNN model and create a model-specific FDR prediction model. Note that, unlike test inputs, training inputs are labeled, and thus mispredicted inputs can be identified. To accomplish this, the entire training set is executed against the original DNN model, the mispredicted inputs are identified, and subsequently, a fault identification approach [18] is applied to these mispredicted inputs. TEASMA thus builds a specific fault prediction model for each individual DNN based on its training set. The performance of TEASMA can potentially be enhanced by integrating new fault identification approaches and creating more precise prediction models.
@@ III-C Sampling Input Subsets
TEASMA selects input subsets from the training set, calculates the AS and FDR of each subset, and builds a regression model. For this purpose, it requires a large number of input subsets that possess diverse AS and FDR values. Therefore, starting with sampling input subsets with a small size, TEASMA updates the sampling size and continues with larger or smaller sizes where needed, until the stopping criterion based on the minimum ($minFDR$) and maximum ($maxFDR$) FDR values achieved by the samples in $A$. To build an accurate prediction model, TEASMA requires the sampling of multiple subsets along the size range. The number of subsets sampled for each size is determined through an input parameter ($SN$) provided to TEASMA, which we set to 300 in our experiments. Sampling is performed with replacement, allowing for the potential inclusion of the same test input in multiple subset samples. This approach serves the purpose of including diverse test inputs and obtaining significant variance in AS and FDR values across samples.
Since one of the MS formulas we use (Equation (2)) considers the number of images from each class $C_{i}$ in an input subset, TEASMA complements the conventional approach of creating randomly sampled subsets by creating a set of uniformly sampled subsets. Random subsets may contain any number of input images from each class $C_{i}$ while uniform subsets contain an equal number for each class $C_{i}$. Uniform subsets may achieve higher MS values than random subsets based on Equation (2). Therefore, we perform our experiments using MS with both types of input subsets and empirically determine which one serves our purpose better.
@@ III-D Regression Analysis
In this step, TEASMA uses all the sampled input subsets from the training set to perform regression analysis between AS and FDR. Given that post-training MOs may not be representative of real faults, we can expect non-linear relationships. In fact, as our subsequent results demonstrate, the shape of the relationship between AS (calculated using MS, SC, and IDC) and FDR is often highly non-linear and significantly varies across DNNs. Therefore, TEASMA considers quadratic and exponential regression models, as well as regression trees when no regression function fits well. Regression trees enable the optimal partition of the AS range into optimal sub-ranges with similar FDR values while avoiding overfitting. Note that the best model shapes may differ across DNNs because of differences in their architecture and training sets. Consequently, for each DNN model, the best shape for the regression model is determined empirically, prioritizing simpler linear models when they fare well. Regardless of its shape, the regression model must be sufficiently accurate to enable practically useful FDR predictions. To measure the goodness of fit for each regression model, TEASMA calculates the resulting coefficient of determination ($R^{2}$). TEASMA also computes two common measures for evaluating the accuracy of predictions, namely the Mean Magnitude of Relative Error (MMRE) and the Root Mean Square Error (RMSE) to assess how accurately FDR can be predicted. To obtain more realistic results, TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average $R^{2}$, MMRE, and RMSE across all folds. Based on these metrics, engineers can determine for each DNN, the best regression model for predicting FDR based on AS.
It is important to note that according to TEASMA’s methodology, a specific prediction model needs to be built for each DNN model using a selected adequacy metric. This is, in practice, not an issue as one relies on the specific training set of each model. Thus, despite the generalizability of TEASMA as an approach, it must be reapplied if the DNN model is updated or if another adequacy metric is employed.
@@ III-E Assessing Test Sets
In the last step, when provided with a test set, TEASMA computes its AS using the original DNN and the same adequacy metric and configuration used in previous steps (e.g., previously generated mutants). Subsequently, it predicts the test set’s FDR ($\widehat{FDR}$), including PIs based on the prediction model obtained from the previous steps. Given a certain confidence level (e.g., 95% in our experiments), such intervals specify the probabilistic range of actual FDR values for test sets achieving a specific AS value and thus provide engineers with valuable insights into the uncertainty of FDR predictions. Depending on the context, one can then decide how conservative should decision-making be by considering the lower or higher bounds of the PI. The $\widehat{FDR}$ of a test set, if deemed sufficiently accurate based on its PI, can then be used to assess its capability to detect faults and therefore whether, in a given context, it is adequate to validate the DNN.
@@ IV Experimental Procedure
In this section, we describe the empirical evaluation of TEASMA, including the research questions we address, the subjects on which we conducted our evaluation, the configurations for calculating AS, and a comprehensive description of our experiments.
@@ IV-A Research Questions
Our experiments are designed to answer the following research questions, relying on four test adequacy metrics for AS calculation: MS, DSC, LSC, and IDC:
RQ1: Can an accurate regression model be built to explain FDR as a function of AS, based on any of the existing adequacy metrics?
In this question, we investigate the output of the second step of TEASMA in Section III. We select a large number of input subsets from the training set and calculate their AS using four existing adequacy metrics (i.e., MS, DSC, LSC, and IDC) and their FDR using the faults estimated in the original model, as described in Section III-B. We further break this question down into two sub-questions.
RQ1.1: Which adequacy metrics have a strong correlation with FDR? In this question, we investigate, for each adequacy metric, if the AS of an input subset is significantly correlated with its FDR. While a high correlation between AS and FDR demonstrates that there is a strong relationship between the two, it is also essential to further investigate the shape of their relationship. If the relationship between AS and FDR consistently follows a linear pattern with a slope of 1 for any of the adequacy metrics, then that metric is a surrogate for FDR and can be used directly to assess the adequacy of the test sets. However, our results, as explained in the next section, demonstrate that not only this is not the case for any of the investigated metrics, but the shape of the relationship greatly varies across subjects, thus justifying the next sub-question.
RQ1.2: Based on which adequacy metrics can an accurate regression model be built to predict FDR from AS based on the training set? For adequacy metrics with a strong correlation between AS and FDR, we resort to regression analysis to model the relationship between AS and FDR. Consequently, we investigate if an accurate regression model can be built to predict the FDR of an input subset from its AS, based on the training set. If that is the case, the predicted FDR can serve as a good basis on which to decide whether a test set provides sufficient confidence about the validation of the DNN before its deployment. Though such questions have been investigated using MS for traditional software [16, 12], that has not been the case for DNNs.
RQ2: Which adequacy metrics lead to better FDR predictions on test sets?
In this question, we evaluate the last step of TEASMA in Section III and investigate separately for each selected adequacy metric, if we can accurately predict a test set’s FDR using the regression model built in the previous step on the training set.
RQ3: How accurately can we predict a test set’s FDR? In this question, we compare the results obtained from the last step of TEASMA using selected adequacy metrics and identify which one can be used to most accurately predict a test set’s FDR, with a reasonably small confidence interval.
@@ IV-B Subjects
TABLE I: Information about experimental subjects
ID Input Training Test Model Epochs Train Test
set set set accuracy accuracy
S1 MNIST 60,000 10,000 LeNet-5 12 96% 96%
S2 Cifar-10 50,000 10,000 Conv-8 20 94% 85%
S3 ResNet20 100 97% 89%
S4 Fashion-MNIST 60,000 10,000 LeNet-4 20 94% 90%
S5 SVHN 73257 26032 LeNet-5 20 94% 87%
S6 VGG16 10 97% 93%
S7 Cifar-100 50,000 10,000 ResNet152 10 89% 79%
S8 ImageNet 1,281,167 50,000 InceptionV3 pre-trained 89% 77%
@@
We performed our study on different combinations of input sets and models. Since we rely on a SOTA fault estimation approach for DNNs that is tailored for image inputs, we selected a set of widely-used publicly available image input sets including MNIST [36], Fashion-MNIST [36], Cifar-10 [37], SVHN [38], Cifar-100 [39], and ImageNet [21]. Most of these input sets have also been used for evaluating SOTA mutation testing tools and operators [2, 1, 3, 4], as well as for evaluating adequacy metrics such as SC and IDC. MNIST is a set of images of handwritten digits and Fashion-MNIST contains images of clothes that are associated with fashion and clothing items. Each one of MNIST and Fashion-MNIST represents a 10-class classification problem and consists of 70,000 grayscale images of a standardized 28x28 size. Cifar-10 is another widely-used input set that includes a collection of images from 10 different classes (e.g., cats, dogs, airplanes, cars). SVHN [38], a set of real-world house numbers collected by Google Street View, can be considered similar to MNIST since it includes images representing a sequence of digits from 10 classes of digits. The Cifar-10 and SVHN input sets contain 32x32 cropped colored images. Moreover, we selected two large-scale image input sets, namely Cifar-100 [39] and ImageNet [21], that contain realistic and feature-rich images. The CIFAR-100 input set is a widely used benchmark in the field of computer vision, consisting of 100 classes of images, each containing 600 images. Among these images, 500 are allocated for training, while the remaining 100 are designated for testing, resulting in a total of 10,000 test images and 50,000 training images. Each image in the CIFAR-100 input set is sized 32×32 pixels and is in full color. ImageNet is a widely recognized large-scale input set designed for visual object recognition research in the field of computer vision. It includes over 14 million labeled images across 1,000 object categories [21]. We leverage its most commonly used subset, ImageNet-1k, from the ILSVRC2012 competition with a training set of 1,281,167 images and a validation set of 50,000 images for evaluation. The images in the ImageNet input set are sized 224x224 pixels, featuring a higher resolution compared to other input sets such as CIFAR-100.
In our experiments, we trained seven SOTA DNN models using the above input sets: LeNet-5 [40], LeNet-4 [40], ResNet20 [41], ResNet152 [41], InceptionV3 [42], VGG16 [29], and an 8-layer Convolutional Neural Network (Conv-8). Since not all of the post-training MOs are applicable to every DNN model, as described in Section II-A1, we selected a diverse range of DNN models with different internal architectures to ensure that all of DeepMutation’s MOs are examined in our experiment as extensively as possible. We should note that we use a pre-trained InceptionV3 model [42] for the ImageNet input set. Additionally, following standard transfer learning practices, we employ a pre-trained ResNet152 model initially trained on the ImageNet input set, and then fine-tune it for the CIFAR-100 input set rather than training it from scratch [43]. This is motivated by the need for computational resource efficiency and improved performance.
@@
All the feasible combinations of DNN models and input sets we experiment with are referred to as subjects. Overall, we perform our experiment on eight subjects including a wide range of diverse image inputs and DNN architectures. Table I lists the details of each subject including the size of test and training input sets, the number of epochs used to train the model, and the accuracy of the DNN model on the training and test input sets.
It is important to note that our experiments were conducted using the original test sets of the subject models, as defined in their respective original sources [36, 38, 37, 39, 21], to eliminate potential biases, as we were not involved in their selection process. The subjects’ DNN models cover a wide but high accuracy range, from 77% to 96%, as detailed in Table I. As a result, scenarios where the model’s accuracy is poor, such as when a distribution shift occurs between the training and test set, are not considered in our experiments. Such scenarios reveal the need for retraining, and thus a test adequacy assessment becomes superfluous, as the goal of such assessment is to provide confidence in high-accuracy test results.
IV-C AS calculation Configuration
Figure 3: Configuring MS, SC, and IDC calculation in our experiments
In this section, we describe the configurations used in our experiments to calculate AS based on alternative adequacy metrics, as illustrated in Figure 3. To calculate MS precisely, we need to complement the mutant generation process by identifying and excluding mutants that are trivial or equivalent. In our experiments with MS, we use DeepMutation++ [3], the most recent SOTA mutation testing tool for DNNs that implements eight post-training MOs suitable for FNNs. Some of the implemented MOs need to be configured by setting values for their parameters. For example, the Neuron Effect Blocking (NEB) operator is configured by determining the number of neurons it affects. By default, the DeepMutation++ tool configures this operator to sample a predefined ratio of 1% of the total neurons in a DNN model and limits the total number of generated mutants to 50. We use the default configuration for generating mutants. Furthermore, we perform the same procedure proposed by the DeepMutation++ tool and use its default thresholds to filter out some of the generated mutants. Specifically, we exclude mutants when they yield less than 90% of the original model’s accuracy or mispredict more than 20% of the inputs correctly predicted by the original model, as reported in the original papers [2, 3]. Additionally, we perform an extra step to filter out equivalent mutants. For this purpose, the entire training set is executed against the mutants and those that cannot be killed by any training input are considered equivalent and therefore excluded. Various approaches can be employed for MS calculation as explained in Section II-A3. Consequently, in our experiment, we examine the correlation between FDR and MS, calculated using each approach, and empirically determine which approach results in the most accurate prediction model. The post-training MOs utilized in our experiments, as implemented in the DeepMutation++ tool [3], are as follows:
  • Gaussian Fuzzing (GF): Alters the weights of a DNN using a Gaussian distribution. This introduces variations in the connection strengths between neurons, subtly modifying the model’s behavior.
  • Weight Shuffling ($WS$): Randomly shuffles the weights of connections for a selected neuron, affecting how inputs are processed by subsequent layers and potentially altering the output.
  • Neuron Effect Blocking ($NEB$): Temporarily disables a neuron’s effect on subsequent layers by setting its outgoing weights to zero. This simulates the impact of a non-functional neuron on the model’s decision process.
  • Neuron Activation Inverse ($NAI$): Inverts the activation status of neurons. By changing the sign of output values before applying activation functions, this operator can alter the behavior of a model.
  • Neuron Switch ($NS$): Swaps two neurons within the same layer, changing their roles and potentially altering the layer’s output pattern.
  • Layer Removal ($LR$): Randomly removes a layer of the DNNs, useful for assessing the redundancy and criticality of particular layers. The $LR$ operator is restricted to layers whose input and output shapes are consistent [2].
  • Layer Addition ($LA$): Adds an additional layer to the model, which can introduce significant changes to the model’s behavior and output. $LA_{m}$ operates under similar conditions as the LR operator to avoid breaking original DNNs.
  • Layer Duplication ($LD$): Inserts a copied layer immediately following its original layer. The LD operates under similar conditions as the LR operator to avoid breaking the original DNNs.
It is important to note that not all MOs are applicable to every DNN architecture. Furthermore, the selection of MOs can affect the effectiveness of mutation analysis and thus, the results of applying TEASMA with MS. Consequently, we investigate the potential impact of MO choices in our experiments. We measure the correlation for mutants generated by each MO separately and then attempt to find subsets of operators leading to higher correlation. We compare the highest correlation achieved by subsets of MOs with that obtained with all MOs. If the correlation of the former is higher than the latter, then we only consider mutants generated by the optimal subset with the highest correlation. Detailed information regarding the applied MOs, on each subject DNN model and their respective configurations, are provided in our replication package [44].
To calculate SC, we use both the LSA and DSA metrics, consequently calculating DSC and LSC scores. When calculating DSA, first, the activation traces of training inputs for a chosen layer in the DNN are captured to establish a reference set. Then, for a given test input, the activation trace is computed at the same layer. Subsequently, the Euclidean distance between the activation trace of the test input and its nearest neighbor among the training traces is calculated, serving as the test input’s DSA score. When calculating DSA for a training input, as required in Algorithm 1, it is necessary to exclude this particular input from the training set; otherwise, the DSA value would be zero. Similarly, to accurately compute LSA for a training input, it is removed from the training set, and the probability density is recalculated based on the remaining inputs. Subsequently, LSA is calculated using the updated density estimation. This process is repeated for each training input to calculate its DSA and LSA values. In our experiments, following the approach of the original paper [6], we calculated DSA and LSA using the deepest (i.e., the last hidden) layer of the model. Additionally, we set the number of buckets for calculating LSC and DSC to 1,000, consistent with the value employed in the original paper [6]. The lower bound and upper bound for calculating the SC of both the training and test subsets are determined using the SA values observed during the training. The details of the DSC and LSC calculation configuration used for each subject are provided in our replication package [44].
The IDC framework instantiation involves training a VAE component, selecting an existing OOD component, and configuring IDC parameters [5]. In the IDC workflow, a VAE is trained with the training set of each input set to model the input distribution in a latent space. In our experiments, we employ the publicly available VAEs from Dola et al. [5] for our first four subjects, which share identical input sets with those used in the original evaluation of the IDC framework. For the remaining input sets, we trained new VAEs. Consequently, in our experiments, six VAEs are employed for the MNIST, Cifar-10, Fashion-MNIST, SVHN, Cifar-100, and ImageNet input sets. All these VAEs use $\beta$-TCVAE [45] model based on the architecture proposed by Burgess et al. [46], with latent sizes of 6, 32, 6, 64, 32, and 64, respectively.
To calculate the IDC of a test set, its test inputs are encoded into the latent space using the VAE’s encoder component. Then test set’s coverage is measured based on the adequacy of its test inputs in covering the partitions of the latent space. Crucially, an OOD detector is employed in the IDC framework to filter OOD test inputs and ensure that only realistic, in-distribution inputs are processed by the VAE’s encoder, making this the most computationally expensive step of the workflow. However, in our experiments, where IDC is computed for test inputs directly selected from the original test set of an input set, the OOD filter step becomes redundant as these inputs are by default in-distribution. Consequently, we exclude the OOD detection step of the IDC workflow in our experiments. Additionally, we use the default IDC parameter configurations from the original implementation of the IDC framework [5]. As shown in Figure 3, calculating IDC for input subsets does not require the presence of the DNN model, making it independent of the model and relying solely on the feature interactions of its inputs.
IV-D Experiments